### v4

# Aligning Perception, Reasoning, Modeling and Interaction: A Survey on Physical AI

Abstract
The convergence of embodied intelligence and world models has catalyzed growing interest in integrating physical laws into AI systems. While prior surveys have examined world models and embodied intelligence separately, we focus on the progression that connects these capabilities as a unified developmental pathway from passive observation to active physical comprehension. This survey provides a systematic framework revealing how physical AI advances through four interconnected stages: perception transforms sensory data into structured physical representations, reasoning derives explanations from observed phenomena, modeling enables predictive simulation grounded in physical principles, and embodied interaction closes the loop through physical manipulation and environmental feedback. Each stage enables and enhances the next: perceptual grounding supports causal reasoning, reasoning unlocks predictive capabilities, and robust models drive genuine physical interaction. Through analysis of developments spanning architectural innovations, training methodologies, causal inference, and embodied systems, we synthesize how physical understanding emerges through cumulative integration across this progression. Our framework reveals the evolution from isolated, task-specific solutions toward integrated architectures that advance from pattern recognition toward causal reasoning and counterfactual prediction. This perspective provides foundations for next-generation physical AI systems with direct implications for safe, generalizable, and interpretable deployment across robotics, scientific discovery, and autonomous systems. We maintain a continuously updated taxonomy repository at https://github.com/AI4Phys/Awesome-AI-for-Physics.
Index Terms:
Physical AI System, Physical Perception, Physics Reasoning, World Modeling, Embodied Interaction.
“What I cannot create, I do not understand.”
Richard Feynman [^1] Richard Feynman (1918-1988) was a Nobel Prize-winning physicist known for his groundbreaking work in quantum electrodynamics and the invention of Feynman diagrams.
A Introduction
Teaching artificial intelligence to understand our physical world represents one of the most fundamental challenges in modern AI research [World_models2018, AIMeetPhy_survey_2024]. While humans naturally grasp complex physical interactions from early childhood, frontier models struggle with basic physical reasoning that young children master effortlessly [NewtonianIU, Battaglia2016InteractionNF]. This capability gap becomes increasingly critical as AI systems are deployed in real-world scenarios ranging from self-driving vehicles to humanoid robots. To quantify this limitation, evaluation frameworks such as SeePhys [seephys] for symbolic reasoning, PHYRE [I-PHYRE] for intuitive physical perception, and PhyBlock [Ma2025PhyBlockAP] for realistic dynamics prediction have emerged, consistently demonstrating that current models lack structured understanding of physical laws. This gap manifests dramatically in performance disparities. A vision model trained on millions of images may achieve superhuman image classification accuracy yet fail to predict elementary outcomes such as a bouncing ball’s trajectory [LLMPhy_2025, ComPhy]. This paradox reveals that contemporary AI models learn statistical correlations from massive data rather than developing causal knowledge rooted in physical principles.
Figure 1: Overview of four physical understanding capabilities of current AI systems. Physical Perception denotes the model’s ability to recognize the properties of objects. Physics Reasoning refers to the model’s capability to understand symbolic systems and physical laws. World Modeling represents its capacity for simulation and prediction. Embodied Interaction characterizes the model’s ability to interact with the physical environment.
Addressing this fundamental challenge requires physics-aware reasoning models across multiple complementary directions. From an architectural perspective, Graph Neural Networks (GNNs) [gcn, gat, graphsage] leverage their inherent capacity to model relational structures and capture pairwise physical interactions between objects [Battaglia2016InteractionNF, VIN, NPE], naturally encoding spatial relationships and interaction dynamics. Diffusion-based models [MotionCraft, VideoREPA] have demonstrated substantial potential in modeling complex physical processes through learned denoising procedures, while transformer architectures [Kantamneni2024HowDTA, Han2025APMA] increasingly incorporate physics-aware attention mechanisms and constraints. Beyond architecture, training methodology plays a crucial role. Researchers have developed physics-guided loss functions that explicitly incorporate physical constraints [elhasadi2019_semisupervised_sr_drag], curriculum learning strategies that progressively introduce physical complexity [chen2025mint], and reinforcement learning frameworks with physics-informed reward structures [AlphaDrive, AutoDrive, liu2025reinforcementlearningmeetslarge]. These training paradigms aim to instill physical intuition during learning rather than relying solely on pattern recognition. At test-time inference, approaches explicitly integrate physical laws and symbolic reasoning into causal modeling [causalmodeling, causalthreads, PhysORD2024], while differentiable physics engines enable end-to-end optimization [differentiable-stokes-flow2020, Scalable-differentiable-physic2020], thereby bridging the gap between learned representations and established physical principles.
Among these directions, large language models (LLMs) and multimodal large language models (MLLMs) offer a particularly promising pathway forward. These models have demonstrated substantial capability in learning from massive datasets and performing sophisticated end-to-end reasoning across multiple modalities [gpt4o, llava, o1, Qwen2.5-VL, Claude-3.7-Sonnet, Gemini-pro]. Multi-agent systems with tool integration have achieved performance comparable to top human competitors in physics competitions [Liang2025MultimodalRF], suggesting potential for more general physical understanding. Generative systems [GAIA-1, DriveDreamer] have achieved controllable, photorealistic synthesis of physical scenarios, while vision-language-action models [KimEtAl2024_OpenVLA, BlackEtAl2024_pi0] integrate natural language instructions with continuous physical manipulation, bridging the gap from abstract reasoning to embodied interaction. These developments reflect a broader paradigm shift in physical AI from isolated, task-specific solutions toward integrated architectures (Figure 2) with synergies across domains, extending beyond recognizing input features toward generating and interacting with realistic physical scenarios.
Scope Comparison and Contributions. As summarized in Table I, existing surveys have examined individual dimensions of physical understanding in isolation, addressing perception [Yin2023ASO, Marjieh2023LargeLM, Sapkota2025ObjectDW, Guo2024PhyGraspGR, Sun2025ProbingPC], reasoning [Zhou2025FromPT, Li2025FromS1, Ravishankara2025TheAI, Duan2022ASO, Huang2022TowardsRI], modeling [Khan2025FoundationMD, Barman2025LargePM, Ding2024UnderstandingWO, Liu2025GenerativePA, Zhu2024IsSA, Kong20253DA4, Bergen2019MachineLF, Xie2025From2T, Chen2025ASO, Mai2024FromEM], and interaction [Liu2024AligningCS, Qi2024ShapeLLMU3, Sathyam2025FoundationMF, Feng2025EmbodiedAF, Long2025ASL, Wang2023ASO, Liang2025LargeME, Zheng2024ASO, Wang2025TowardEA, Duan2021ASOA, Xu2024ASOA, Han2023ASOA] as separate research areas without examining the synergistic connections between them. Our survey uniquely focuses on the evolutionary trajectory that unites these four capabilities into a coherent paradigm, analyzing how they interact and inform one another toward unified physical AI systems. We examine how deep learning systems leverage the laws of physics to solve physics problems in an end-to-end manner, as opposed to how physics principles inspire neural network architectures. We leave physics-inspired architectures such as Boltzmann Machines and Hopfield Networks, as well as the broader field of deep learning for physics research, for dedicated AI4Science surveys. To this end, we present a three-tier taxonomy that systematically organizes research across these four fundamental capabilities (Figure 1), with hierarchical task structures and detailed methodological analysis for each. Drawing from over 300 papers, our coverage spans advanced physics problems to applied tasks including object recognition, spatial perception, video generation, robotic control, and autonomous driving, revealing the integrated development pathway from passive observation to active physical comprehension.
Survey Structure. Section B presents the proposed taxonomy. Sections D through F systematically examine each reasoning capability with corresponding tasks and methodological approaches. Section G outlines promising future directions toward more capable physical AI systems. Section H synthesizes our contributions and findings.
Figure 2: Timeline of the development of Physical AI Systems. The timeline illustrates the evolution of key research works, organized into chronological periods. The colored blocks indicate the primary research theme of each paper: pale mint for Physical Perception, soft green for Physics Reasoning, cyan for World Modeling, and sky blue for Embodied Interaction.
TABLE I: Scope of existing surveys. They tend to examine only single dimensions of physical understanding abilities, without providing a comprehensive overview of physical AI evolution.
Capability Existing Surveys
Perception [Yin2023ASO, Marjieh2023LargeLM, Sapkota2025ObjectDW, Guo2024PhyGraspGR, Sun2025ProbingPC]
Reasoning [Zhou2025FromPT, Li2025FromS1, Ravishankara2025TheAI, Duan2022ASO, Huang2022TowardsRI]
Modeling [Khan2025FoundationMD, Barman2025LargePM, Ding2024UnderstandingWO, Liu2025GenerativePA, Zhu2024IsSA, Kong20253DA4, Bergen2019MachineLF, Xie2025From2T, Chen2025ASO, Mai2024FromEM]
Interaction [Liu2024AligningCS, Qi2024ShapeLLMU3, Sathyam2025FoundationMF, Feng2025EmbodiedAF, Long2025ASL, Wang2023ASO, Liang2025LargeME, Zheng2024ASO, Wang2025TowardEA, Duan2021ASOA, Xu2024ASOA, Han2023ASOA]
B Taxonomy
Physical AI Systems Physical Perception Physics Reasoning World Modeling Embodied Interaction Object Recognition Category & Scene Recognition Anomaly Detection Spatial Perception 2D & 3D Relationships Metric-Semantic Mapping Identifying Intrinsic Property Material & Optical Properties Mass & Rigidity Estimation Dynamic Estimation Interaction Dynamics State Prediction Causal and Counterfactual Reasoning Causal Discovery Intervention-based Learning Benchmarking Physics Reasoning Text-based Benchmarks Multimodal Benchmarks General Reasoning Models Research-Level AI Scientist Uni-Level PreUni-Level Theoretical Physics Solvers Symbolic Regression Constraint-based SR Physics-Informed Networks (PINNs) PINN Applications Generative World Modeling Image Generation Video Generation Scene Reconstruction Physics-enhanced World Modeling Neuro-symbolic Integration Physics-Structured Neural ODEs Differentiable Physics Engines Benchmarking World Modeling Generation Quality Evaluation Dynamic Modeling Evaluation Virtual Interaction Capability Robotics Continuous Action Generation Cross Platform Generalization Real-world Perceptual Grounding Navigation Object-Goal Navigation Vision-and-Language Navigation Dialog-based Navigation Autonomous Driving Rule-based Approaches Learning-based Approaches World Models & Generative Hybrid Approaches
Figure 3: The proposed taxonomy of AI systems for physics understanding capabilities, organized into four fundamental capabilities. Each capability is further broken down into hierarchical sub-topics that reflect an increasing level of cognitive complexity or domain specificity.
To systematically understand how AI systems engage with physics, we propose a hierarchical taxonomy that reflects the natural progression from abstract reasoning to embodied action. Our framework organizes physics-aware AI capabilities into four interconnected domains that mirror human cognitive development in physical understanding. Physical Perception represents the ability to intuitively extract physical properties from sensory data through increasingly sophisticated layers, including object recognition, spatial relationships and intrinsic properties. In the inverse direction, Physics Reasoning encompasses symbolic manipulation and theoretical problem-solving, where AI systems leverage mathematical representations to solve physics problems ranging from textbook exercises to research-level challenges. These two complementary pathways converge in World Modeling, where AI systems integrate symbolic knowledge with perceptual understanding to construct predictive models of physical environments, enabling capabilities from video generation to reconstruction. Finally, Embodied Interaction grounds these cognitive capabilities in real-world action, where robots, autonomous vehicles, and navigation agents must reconcile theoretical understanding with the irreversible consequences of physical manipulation. Figure 3 shows our taxonomy. This taxonomy not only delineates current research boundaries but also reveals critical gaps where symbolic reasoning and embodied experience remain disconnected, highlighting the path toward genuinely physics-aware artificial intelligence.
C Physical Perception: From Sensory Data to Physical Understanding
Understanding the physical world begins with perception, where AI systems transform raw sensory inputs into structured representations of physical properties, spatial relationships, dynamic behaviors. Physical perception serves as the essential grounding for subsequent reasoning, modeling, embodied interaction, establishing the observational foundation upon which symbolic reasoning and predictive models are built. While traditional computer vision focuses on semantic recognition, physical perception demands extracting geometry, material properties, force dynamics, causal structure from visual or multimodal observations. This capability enables systems to infer latent physical states that are not directly observable but fundamentally govern real-world phenomena. We organize existing investigations along a progression of increasing cognitive complexity: Object Recognition extracts geometric structure with semantic information, Spatial Perception captures relationships within scene composition, Intrinsic Property Estimation infers material characteristics, Dynamic Estimation predicts motion through temporal evolution, Causal and Counterfactual Inference reveals underlying physical mechanisms. This hierarchical organization reflects how perceptual capabilities build cumulatively, with each level providing foundations for the next, ultimately enabling the transition from observation to reasoning about physical laws.
C.1 Object Recognition
The most fundamental aspect of visual physical perception is the ability to identify objects and determine their spatial relationships within a given scene. In the past decade, the development of convolutional neural networks (CNNs) has made it possible to solve target detection and object classification problems. GPT-4V [GPT4V] serves as a milestone that demonstrates robust zero-shot object detection and localization capabilities of MLLMs across diverse visual contexts. Other open-sourced models have also revealed that the introduction of high-quality labeled data can help agents recognize objects at multiple levels of granularity, from basic categories (e.g., "vehicle," "animal") to fine-grained classifications (e.g., "Maserati," "golden retriever") [he2018maskrcnn, liu2024groundingdinomarryingdino, xu2025qwen3omnitechnicalreport]. For more complex scene-level recognition tasks, MLLMs should integrate individual object detections into a coherent understanding of the environment. This involves recognizing scene categories (indoor/outdoor, kitchen/bedroom), understanding typical object arrangements, and identifying anomalous configurations [Song_2024, varghese2025viewinvariantpixelwiseanomalydetection, zhang2024cognition, Ji_2024]. Recent evaluations [MME, Roboflow100-VL, EagleVision, ROD-MLLM] have shown that MLLMs have achieved competitive performance on object-level detection capabilities, which lays the foundation for subsequent physical property perception works.
C.2 Spatial Perception
Beyond object recognition, AI systems must understand spatial relationships to build coherent scene representations. This includes both absolute positioning (e.g., "at the center of the whole image") and relative positioning (e.g., "to the left of the white table"). Many recent benchmarks are specifically designed to test spatial understanding, such as VSR and SpatialBench [SpatialScore, Open3DVQA, yang2025mmsi, wu2025spatialmllmboostingmllmcapabilities, xu2025multi], reveal varying capabilities across different MLLMs. Models generally perform well on basic 2D spatial prepositions (above, below, left, right) but struggle with more complex spatial concepts such as 3D spatial reasoning, pixel-level localization and scale relationships. The ability to handle these complex spatial tasks is typically limited by the nature of their training data and architectures. Most models rely on large-scale datasets like COCO [coco] or ImageNet [Imagenet], which focus on object identification and simple relationships but lack detailed annotations for more sophisticated spatial reasoning tasks. To overcome these limitations, recent research has begun to explore incorporation of spatially aware neural networks [SpIRL, SpaGAN, benou2025tellvisuallyexplainabledeep, cho2024spatially] and multi-view learning [yu2025review, ViewSpatial, zheng2023comprehensive, zhang2024multi] to handle complex spatial relationships.
C.3 Identifying Intrinsic Property
Understanding the physical world from vision requires not only recognizing objects but also inferring their intrinsic properties and dynamic behaviors based on these properties. Intrinsic properties such as mass, viscosity and rigidity are inherent characteristics of objects that remain constant regardless of observational perspectives. Estimating these properties from visual observation alone is particularly challenging for AI as it demands mapping visual features to physical attributes that may not be directly observable but can only be inferred by the laws of physics. Recent studies aim to achieve reliable identification of material such as metal, fabric and plastic [kocsis2024intrinsicimagediffusionindoor, Van_Zuijlen_2021] and even finer textures (e.g., silk vs. cotton) [hu2020fabricsurfacecharacterizationassessment, jang2024fabric], while also identifying optical properties such as transparency and translucency [chen2024practical, weidenbach2024transparency, nagai2025top]. Nevertheless, correctly identifying ambiguous cases, novel materials, and lighting variability remain major challenges for frontier models. For mass and weight estimation, models primarily rely on size cues, material-density associations, and category priors, showing stronger performance in relative rather than absolute weight prediction, though systematic biases persist [Vision-Based, afridi2024analyzing, kalus2024simulating, M_ller_2024, nath2024mass]. In terms of rigidity and deformability, MLLMs can classify rigid and flexible objects, detect evidence of deformation, and exploit contextual cues such as force interactions, but their judgments remain heuristic and fragile in complex settings [PhysID, mustafa2025forcemapping, song2025moda, PaNDaS].
C.4 Dynamic Estimation
Crucially, these intrinsic properties serve as foundation for understanding how objects behave according to the laws of physics. Knowledge of mass and rigidity, for instance, directly informs whether/how an object might deform under external force, or how it influences motion during collisions. Building on this static perspective, dynamic property perception captures how objects behave and interact over time through contact, constraints, and forces such as support, occlusion, friction, and impact. Unlike intrinsic estimation that answers what an object is in terms of characteristics, dynamic perception addresses how it behaves upon interaction with other objects. Graph-based methods, especially graph neural networks (GNNs), have been central to this effort, with models such as Interaction Networks, Visual Interaction Networks, and the Neural Physics Engine [VIN, NPE] demonstrating how object relations can be captured to predict physical dynamics from visual inputs. More recent frameworks like I-PHYRE [I-PHYRE] further challenge agents to exhibit intuitive physical reasoning, multi-step planning, and in-situ intervention tasks, highlighting the importance of real-time dynamic perception. In addition, benchmarks such as DeepPHY [DeepPHY] provide systematic evaluation environments that couple intrinsic attributes with dynamic outcomes, ensuring models are tested on both the static and temporal dimensions of physical understanding. By moving from the recognition of intrinsic object properties to the perception of their dynamic relations, AI systems develop a more comprehensive human-like understanding of physical environments.
C.5 Causal and Counterfactual Inference
Beyond perceiving object dynamics from massive aggregated data, causal and counterfactual inference represents the most sophisticated layer of physical reasoning, which focus on elucidating how underlying mechanisms that govern observed phenomena enable reasoning about hypothetical "what if" scenarios without direct observation. Rather than merely describing temporal event sequences, causal reasoning addresses more fundamental questions such as why events occur and how modifications to underlying conditions would alter seemingly irrelevant outcomes. For instance, altering gravitational constants would fundamentally change the falling dynamics of objects, exemplifying the advanced reasoning capabilities required to address such counterfactual queries that transcend pattern matching from observational data. Contemporary methodologies in this domain encompass causal graph modeling [causalmodeling], intervention-based learning frameworks [ke2020causalmodels], and hybrid architectures that integrate neural networks with physics engines or symbolic reasoning systems, as demonstrated in PhysORD [PhysORD2024]. Evaluation of these approaches relies on the development of various physical reasoning benchmarks: the Causal3D benchmark [causal3D] provides structured data with corresponding visual representations; the CLEVRER-Humans dataset [clevrerhumans] offers video-based reasoning tasks for causal judgment of physical events with human annotations, addressing the need for diverse event types and natural language descriptions; and the recent PhySense benchmark [physense] emphasizes principle-based reasoning by systematically evaluating whether large language models can apply human-like physical principles. By capturing latent causal structures, this layer provides interpretable, generalizable, and mechanistically grounded reasoning capabilities, representing the pinnacle of visual physical understanding.
D Physics Reasoning: From Observation to Explanation
Physics reasoning represents the critical transition from perception to explanation, where systems move beyond identifying physical properties to understanding why phenomena occur. Building directly upon perceptual foundations, reasoning operates through equations and theoretical principles that encode physical laws. While perception answers "what is happening" by extracting observable properties, reasoning addresses "why it happens" by connecting abstract concepts with concrete phenomena through causal relationships governed by natural laws. This capability bridges observational understanding with predictive modeling, enabling systems to not merely recognize patterns but derive explanations grounded in physical principles. We examine how AI systems perform physics reasoning across two domains of increasing sophistication: general physics problems spanning textbook exercises to competition-level challenges, and theoretical problems requiring research-level mathematical derivation with conceptual synthesis. This progression reflects the deepening cognitive demands from applying established principles to discovering novel theoretical insights, ultimately preparing systems for the modeling stage where learned physical laws enable forward prediction and counterfactual simulation.
Figure 4: Error pattern analysis of closed-Source models in multimodal physical reasoning.
TABLE II: Performance (%) of LLMs on representative text-based and multimodal physics benchmarks. GPT: GPT-4o [gpt4o], Claude: Claude-3.7-Sonnet [Claude-3.7-Sonnet], DeepSeek: DeepSeek-R1 [guo2025deepseek], Gemini: Gemini-2.5-pro. [Gemini-pro]
Benchmark GPT Claude DeepSeek Gemini
Text-based
UGPhysics [ugphysics] 38.66 56.34
PHYBench [qiu2025phybench] 6.97 12.87 33.45 45.82
GPQA [GPQA] 50.3 67.4 71.5 84.0
OlympiadBench [olympiadbench] 39.72 7.34
Multimodal
SeePhys [seephys] 21.9 34.6 42.2 54.9
PhysReason [physreason] 29.58 34.07
MMMU [mmmu] 59.4 75.0 81.7
MMMU-Pro [mmmupro] 51.9 76.4
D.1 Benchmarking Physics Reasoning
We first systematically review the current landscape of physics reasoning benchmarks from both text-based and multimodal perspectives, shown in Table II.
Text-based Benchmarks. Early benchmarks primarily focused on textbook physics problems in pure textual format, establishing baselines for symbolic manipulation and numerical reasoning.
Representative resources such as PhysicsEval [physicseval] compile problems across multiple education levels and emphasize inference-time techniques, numerical calculation, and robustness under distributional shifts. These benchmarks provide systematic coverage from fundamental mechanics and kinematics to electricity, thermodynamics, and modern physics, enabling evaluation of both single-step calculations and multi-step reasoning chains. Other text-based datasets, including UGPhysics [ugphysics], PHYBench [qiu2025phybench], GPQA [GPQA], further extend the scope of questions to university-level courses and examinations.
Empirical studies consistently show that even state-of-the-art language models often depend on superficial pattern matching and memorization, with sharp performance degradation on tasks requiring genuine physical reasoning, long multi-step inference, or generalization beyond the training distribution. Such findings highlight the persistent gap between current models and the robust reasoning capabilities expected from human-level or world-model-based physical understanding, motivating future benchmarks that more explicitly target multi-step reasoning, causal inference, and compositional problem-solving.
Multimodal Benchmarks. On another dimension, a fundamental distinction between physics reasoning and pure mathematics lies in the involvement of more complex visual perception and diagram comprehension, which has driven the development of multimodal benchmarks.
Typical multimodel benchmarks like SeePhys [seephys] highlight the importance of diagram-dependent physical reasoning by organizing problems across seven domains and twenty-one categories, demonstrating that visual context is often indispensable for accurate solutions. Building on this line, PhysReason [physreason] expands the paradigm by introducing multimodal diagnostic tasks and stepwise scoring to assess models’ ability to integrate perceptual cues with symbolic reasoning. More recent benchmarks, including ScienceQA [scienceqa] for science-oriented diagram comprehension, MME-SCI and MMMU series [mme_sci, mmmu, mmmupro] for large-scale perception–reasoning evaluation, and SciEval [scieval] for competition-level and interdisciplinary multimodal science problems, further broden the evaluation landscape by emphasizing visual perception, causal inference, and counterfactual reasoning under dynamically varied conditions. Together, these benchmarks extend the evaluation focus from purely text-based problem solving to scenarios where perception and reasoning must jointly operate in complex visual environments.
As shown in Figure 4, performance analyses on these multimodal benchmarks consistently reveal that current models excel at pattern recognition but struggle to connect perceptual input with underlying physical principles, especially in long-horizon or distribution-shifted scenarios. This gap underscores that progress in physical reasoning will require not only more sophisticated perception modules but also architectures capable of structured causal reasoning and compositional inference. Developing benchmarks with richer multimodal tasks and more realistic dynamic scenes will be crucial for promoting models capable of integrating perception, reasoning, and simulation into a coherent world-model of physical systems.
D.2 General Physics Reasoning Models
AI systems for general physical reasoning have evolved from simple programs to sophisticated autonomous agents. These systems must tackle complex problems requiring advanced deduction and abstraction of real-world scenarios. Early studies used attention mechanisms to jointly identify salient information in text and diagrams for scientific QA tasks [GmezPrez2020ISAAQMC]. More recent work in Prompt Engineering and Mixture of Experts (MoE) uses CoT prompting to elicit analogical reasoning [Addala2024StepsAAA, Jaiswal2024ImprovingPR]. Decomposing problems or reasoning processes into simpler atomic components has also been shown to improve reasoning performance, as demonstrated by instruction fine-tuning methods [xiang2024atomthink, kuang2025atomic] and knowledge graph-based approaches [Addala2024KnowledgeGAE].
A significant leap is the development of agentic systems. A key example is The AI Scientist, which generates research papers from ideation to final manuscript [lu2024aiscientist]. Its successor, The AI Scientist-V2, was the first AI system to have a paper accepted through peer review [yamada2024aiscientistv2]. Other agents demonstrate performance equal to human experts in high-level competitions. In particular, Physics Supernova [PhysicsSupernova] ranked 14th out of 406 contestants at the 2025 International Physics Olympiad (IPhO).
Multi-agent systems have also shown great value for scientific discovery [CollectiveIntelligence] as we demonstrate in Figure 5. For instance, AtomAgents helps design novel metal alloys [ghafarollahi2024atomagents]. The Multimodal Universe project processes vast astronomical datasets and has compiled over 100 TB of data for machine learning research [qu2024multimodal]. Visual and multimodal reasoning remains a critical challenge. One successful approach used text descriptions of images to help AI better understand physics diagrams, which won first place at the ICML 2025 AI4Math Challenge [Liang2025MultimodalRF]. These advancements show a clear trajectory toward automating scientific workflows. However, the field still relies on specialized systems, particularly for tasks with strong visual dependence.
Figure 5: Multi-agent physics reasoning system across 3 use case spanning theoretical, computational and experimental research.
D.3 Theoretical Physics Solvers
The synergy between artificial intelligence and the physical sciences has evolved into two primary paradigms [carleo2019_ml_physics_survey, AIMeetPhy_survey_2024]: data-driven scientific discovery and physics-informed learning. These approaches provide the methodological foundation for the advanced reasoning systems discussed previously, each addressing the challenge of integrating physical principles in a distinct way.
Data-driven approaches aim to discover unknown physical laws directly from data. A prominent example is Symbolic Regression (SR), which extracts interpretable mathematical expressions without prior knowledge of the governing equations. This makes SR a powerful tool for generating human-readable models that overcome the "black-box" nature of many neural networks. The primary obstacle for SR is navigating its vast combinatorial search space while ensuring physical plausibility. To this end, recent advancements focus on incorporating hard physical constraints, such as unit consistency, to accelerate the learning process and guide the discovery toward physically meaningful equations [zhang2024_interpretable_turbulence_sr].
On the other hand, physics-informed learning is applied to problems where the governing equations are already established. This paradigm incorporates known physical laws as constraints within neural architectures. The most notable example is Physics-Informed Neural Networks (PINNs), which embed the governing partial differential equations (PDEs) directly into the neural network’s loss function. This ensures that the model’s solution adheres to fundamental physical principles. This methodology has found widespread application across diverse fields, including the modeling of high-speed and hidden fluid mechanics [mao2020_pinn_high_speed_flow, raissi2020_hidden_fluid_mechanics], solving the incompressible Navier-Stokes equations [jin2021_nsfnets], and simulating flow in porous media [almajid2022_pinn_porous_media]. Its utility also extends to solid mechanics for problems in elasticity theory and fatigue life prediction [vahab2021_pinn_biharmonic_elasticity, zhou2023_pinn_fatigue_life], as well as to energy systems for solving optimal power flow problems and modeling the complex degradation processes in lithium-ion batteries [nellikkath2022_pinn_ac_opf, wang2024_pinn_battery_degradation].
E World Modeling
The previous sections have traced two key paths, namely Physical Perception and Physics Reasoning, towards AI systems that perceive and understand the law of physics. Physical perception embodies the end-to-end learning approach, extracting physical patterns and regularities directly from sensory observations, while physics reasoning encodes scientific knowledge into symbolic representations and theoretical frameworks.
After internalizing such physical knowledge, can models strengthen their understanding of governing principles of the world and thereby predict the future? World models, as an emerging line of research, offer a natural next step for physical AI systems by enabling a progression from understanding to modeling. World Modeling represents the capability of AI systems to generate, understand and predict the dynamics of simulated environments, enabling a wide range of applications from video generation and scene reconstruction to autonomous planning and decision-making.
Models with world modeling ability have three fundamental advantages for current AI development: (1) reducing the need for supervised data, (2) closing the simulation-to-reality gap, and (3) enabling low-cost, interpretable prediction of future states. These advantages arise because world models enable agents to learn through interaction with simulated environments rather than relying on massive amounts of data, and they can provide more fine-grained sensory data for embodied agents [understanding_or_pridiction].
Unlike physics reasoning LLMs, world models exhibit substantial shifts in their architectural design, training paradigms, and evaluation methodologies. In this section, we categorize litertures of world models into three dimensions: (1) Generative Models: They rely on gradient descent algorithm to learn underlying dynamics of the world from large scale database; (2) Physics-enhanced Approaches: They are constructed based on the principles and laws of physics in symbolic representation, combining the principle of physics that govern our physical world with neural networks to simulate changes in world states; (3) Benchmarks: They provide the metrics used to evaluate model’s capabilities in qualitative and quantitative estimation for world modeling.
E.1 Generative Models
Image Generation. Image generation models transform abstract information into static visual representations. This process encompasses not only traditional lighting models like Phong Shading [Phong] and material rendering like Cook-Torrance [Cook-Torrance], but more crucially utilizes Physically Based Rendering (PBR) through physical simulators like NVIDIA Isaac Sim [isaacsim] and Unity ML-Agents [Unity-ML-Agents] to simulate the interactions between light rays and objects, thereby producing images with high fidelity in both visual and physical perspectives. Recent work in computer vision has demonstrated the effectiveness of PBR towards closing the sim-to-real gap, with applications ranging from depth sensor simulation  [stereovision-depth-sensors] to 6D object pose estimation  [BOP], where PBR-generated training data significantly outperformed traditional rendering approaches.
Under the world model framework, image synthesis establish the connection between internal physical state representations and external visual observations. Contemporary neural rendering approaches, especially those based on Neural Radiance Fields (NeRF), such as Ref-NeRF [Ref-NeRF] and ENVIDR [ENVIDR], allow the rendering process to produce both visually convincing images and physically consistent results. This transformation process from physical simulation to image synthesis effectively translates the internal state representations of world models into an observable and interpretable form for embodied agents, establishing a critical visual foundation for temporal modeling and dynamic prediction.
Video Generation. Video generation refers to the computational technique of automatically creating continuous video sequences from input data such as text, images, or noise using algorithmic methods. As we extend from static image synthesis to dynamic video generation, physical constraints become increasingly complex and important.
Figure 6: Video Generations by GPT4Motion (Figure courtesy of [GPT4Motion]).
TABLE III: Performance of Video Generation Models on representative physical and world modeling benchmarks.
Model PhysicsIQ [Physics-IQ]()($\uparrow$) PhyGen [phygen]()($\uparrow$) VideoPhy [2024videophy] $(\uparrow)$ WorldModel Bench[worldmodelbench]()($\uparrow$)
Sora [opensora] 0.10 0.44 0.28 6.11
Pika [Pika] 0.13 0.44 0.29
CogVideoX [CogVideoX] 0.45 0.49 7.31
LaVie [wang2023lavie] 0.36 0.41
Kling [Kling] 0.49 8.82
Modern video generation models, such as diffusion-based approaches like GPT4Motion [GPT4Motion] and VLIPP [VLIPP], are progressively constrained by physics laws, making the generated video sequences not only visually coherent but also physically plausible. For example, as illustrated in Figure 6, GPT4Motion is constrained by physics laws through physical simulators like PyBullet [PyBullet] by using GPT-4 to generate Blender scripts that simulate realistic physics dynamics, ensuring both visual consistency and adherence to physical laws across video frames, while VLIPP employs chain-of-thought reasoning in vision-language models to predict physically plausible motion trajectories that guide diffusion models to generate temporally coherent and physically accurate video sequences. DINO-Foresight [Karypidis2024DINOForesightLI] supports multiple future understanding tasks and avoids reconstructing irrelevant details of pixel based methods.
This physics-constrained video generation capability is crucial for building reliable world models, as it serves as an evaluation metric of world models’ understanding of environmental dynamics. As shown in Table III, recent video generation models demonstrate varying performance across physical and world modeling benchmarks, indicating their capability to capture environmental dynamics. The ability to generate physically plausible video sequences indicates that the underlying world model has developed robust representations of environmental dynamics, making such generation capability a valuable benchmark for assessing world model quality.
Scene Reconstruction. Scene reconstruction is a computer vision technique that recovers the three-dimensional geometric structure and appearance information of a scene from two-dimensional images, point clouds, or other sensor data, aiming to construct a complete digital 3D representation of the scene. Unlike the forward process discussed earlier that transforms physical states into visual representations, scene reconstruction is an inverse process that requires inferring complete 3D geometry, physical properties, and scene dynamics from limited observational information.
Traditional 3D reconstruction methods are typically based on geometric methods like SFM [SFM] and MVS [MVS], so they demonstrate strong geometric perception capabilities and often neglect the significance of physical reasoning, resulting in reconstructed scenes that violate fundamental physical principles.
Under the world model framework, 3D scene reconstruction should not only recover static geometric structures, but also understand and reconstruct the physical properties and dynamic behaviors among different objects within the scene. This requires reconstruction algorithms to possess not only geometric perception capabilities, but also physical reasoning capabilities. Modern scene reconstruction methods, particularly neural implicit representation-based approaches like PhyRecon [PhyRecon] and IDR [IDR], are making progress in this direction. Not only can they reconstruct precise geometric structures, but also learn and encode physical properties such as density, stiffness, and friction coefficients through physical simulators like MuJoCo [MuJoCo] and Warp [warp2022]. This capability enables reconstructed scenes to be not only visually realistic but also accurate in physical behavior, which is precisely the key requirement for building actionable world models.
E.2 Physics-enhanced Modeling Approaches.
The core challenge facing existing world models is insufficient out-of-distribution physical reasoning capability. While large-scale systems like GAIA-1 [GAIA-1] and DINO-world [DINO-world2025] excel at in-distribution tasks and can generate realistic future scenarios, when encountering counterfactual reasoning, novel object interactions, or unfamiliar physical constraints, these models rely on "case-based" rather than "rule-based" generalization, leading to dramatic performance degradation [physical-law-perspective2024]. Although early physics-based simulation methods can maintain consistency with real-world physics through predefined physics engines [differentiable-stokes-flow2020] [Scalable-differentiable-physic2020], their reliance on manually designed parameterization limits their generalization capability to unmodeled physical effects.
To achieve genuine physical understanding, world models need to accomplish the following critical tasks: (1) Counterfactual physical reasoning: accurately predicting physical behaviors in unseen scenarios; (2) Novel interaction modeling: handling complex inter-object interactions that did not appear during training; (3) Symbolic-grounded knowledge bridging: connecting abstract physical laws with concrete perceptual experiences; (4) Physical constraint generalization: maintaining reasoning accuracy under new constraint conditions. These tasks require systems to maintain strict adherence to fundamental physical principles while adapting to complex and dynamic real-world environments.
To address these challenges, researchers have proposed the concept of ’Neural Physics’, a hybrid framework that preserves fundamental physical laws while using neural networks to learn components that are traditionally difficult to model [neural_physics_2025]. In this way, world models can perform accurate physical reasoning through the combination of physical principles and neural networks. Furthermore, they can create systems capable of bridging symbolic knowledge and embodied experience, representing a hallmark of human-level environmental understanding.
Neural physics methodologies merge explicit physical computations with neural network learning that captures complex material behaviors through data-driven optimization, while preserving end-to-end differentiability. Current approaches can be categorized into three main paradigms:
(1) Neuro-symbolic integration directly embeds known physical laws into the neural network architecture to ensure strict compliance with fundamental physical principles, with the neural network solely responsible for learning uncertain components that are challenging to model accurately. For example, PhysORD [PhysORD2024] models vehicles as controlled Lagrangian systems by directly embedding the Euler-Lagrange equations as hard constraints within the network, where the neural network only needs to estimate quantities that are difficult to compute accurately, such as potential energy gradients and external forces. This approach ensures strict enforcement of fundamental physical laws like energy conservation, thereby maintaining good generalization capability even under data-scarce conditions.
(2) Physics-Structured Neural ODEs retains the mathematical structure of physical equations in the neural ODE framework, where neural networks parameterize various components of the physical equations to enable continuous-time domain modeling. For example, MoSim [MoSim2025]decomposes the rigid body dynamics equations into predictor and corrector parts. The predictor strictly follows the mathematical structure of rigid body dynamics, with inertia matrices, gravity terms, and control forces parameterized through specialized neural network modules. The corrector uses standard residual networks to address complex phenomena such as friction and collisions that are challenging to model explicitly. Time integration is performed via neural ODE solvers, enabling the system to maintain both the inherent structure of physical equations and the learning capacity of neural networks.
(3) Differentiable Physics Engines employs a decoupled architecture in which neural networks specialize in predicting physical properties and interaction parameters from sensor data, with these predictions subsequently input into differentiable classical physics solvers for final system state computation. FusionForce [fusionforce2025] utilizes deep neural networks to predict terrain physical properties such as geometric shape, friction coefficients, and stiffness, along with robot-terrain contact forces from camera images and LiDAR data. These predictions are subsequently fed into a differentiable rigid body dynamics solver for robot motion trajectory computation. The end-to-end differentiable framework allows the model to refine terrain property predictions by backpropagating trajectory errors.
These methodologies demonstrate how physical reasoning emerges from the synergistic interaction between neural learning and constraints offered by the law of physics, moving beyond pattern memorization toward principled understanding of environmental dynamics.
E.3 Benchmarking Modeling Capability
Theperformance of unified world models is best demonstrated through their applications in challenging domains that require genuine physical understanding rather than pattern matching.
Generation Quality Evaluation. Generation quality evaluation assesses the fidelity and coherence of world models’ predictive outputs across multiple modalities. Generative quality evaluation can be mainly categorized into four primary types:
(1) Fidelity evaluation focuses on the degree of similarity between model-generated content and real-world data. Image generation employs metrics such as FID [FID] and CLIP Score [Clipscore], with evaluation conducted on datasets including COCO-30K [coco] and ImageNet [Imagenet]. Video generation usually uses feature methods like I3D [I3D] , with evaluation conducted on Kinetics-400 [Kinetics-400].
(2) Consistency evaluation examines the logical coherence of generated content under different conditions. This includes identity preservation tests on CelebA-HQ [CelebA-HQ], action consistency tests on UCF-101 [Ucf101], and object state change evaluation on the Something-Something dataset [Something-Something].
(3) Spatio-temporal coordination evaluation targets the spatial-temporal coordination capability of video generation. Commonly used benchmarks include Moving MNIST [MovingMNIST] motion prediction, BAIR Robot Pushing [BAIR] object interaction evaluation, and Kinetics-700 action-scene coordination tests.
(4) Physical plausibility assessment evaluates whether generated content adheres to physical laws, utilizing NYU Depth V2 [NYU-Depth-V2] for image depth consistency tests, CLEVR-Video [CLEVR] for video physical reasoning tests, and the Real World Physics dataset for assessing gravity and collision prediction accuracy.
Dynamic Modeling Evaluation. Evaluating the dynamics modeling capability of world models is crucial for assessing their understanding of physical laws and temporal evolution. Current evaluation frameworks can be classified by multiple dimensions: (1) general trajectory prediction: For moving objects, models must forecast future positions based on initial conditions and physical constraints [Bullinger20173DTRA, PDDL, Zhang2024M3BenchBW, Zhu2023BenchmarkingSV]; (2) deformable material simulation fidelity: Require modeling of elasticity and collision dynamics, particularly for complex materials like cloth and rubber [Wang2018NeuralML, Liang2019DifferentiableCSA, Mianroodi2021LosslessMCA]; and (3) fluid dynamics modeling: Test model ability to capture turbulence, viscosity, and flow patterns [predictivemodels, Zhao_2024, liquidsgraphnetworks]. Some comprehensive benchmarks also provides standardized evaluation scenarios covering rigid-body collisions, soft-body deformations, and projectile motion, and is widely used to measure models’ short-term prediction accuracy and long-term stability [Physion, PhysionPlus, ContPhy, PhyWorldBench]. Critical metrics in these benchmarks include physical consistency (adherence to conservation laws), state transition accuracy (alignment between predicted and ground-truth trajectories), and generalization to novel physical configurations.
Virtual Interaction Capability. The ability to accurately model virtual interactions is fundamental to evaluating world models’ understanding of cause-and-effect relationships and their capacity for counterfactual reasoning. Virtual environments like game engines provide an ideal crucible for testing the predictive accuracy and planning capabilities of an agent’s internal world model.The AIBIRDS competition exemplifies this approach, requiring agents to solve unseen levels in real-time through accurate simulation of complex physical interactions [AIBIRDS], thus testing both predictive accuracy and adaptive planning capabilities within physically embedded systems [physicallyembeddedplanningproblems]. Some standard benchmarks such as Hi-Phy, PHYRE and PhyQ assess generalization across diverse interaction templates [Hi-Phy, phyre, Phy-Q], while adversarial evaluations probe deeper understanding through deceptive scenarios that punish superficial heuristics and demand sophisticated multi-step reasoning [deceptive]. Other works also demand that explicit physics simulators deliver interpretable, physics-aware decision support in complex interaction scenarios [CueTip]. Ultimately, these virtual evaluating environments demonstrate that achieving robust physics intelligence requires more than just prediction, but involving human-like decision-making strategies such as knowing when to stop loss and restart a failed attempt [usingrestartheuristicsimprove], where this kind of trial-and-error learning fundamental to solving real-world challenges like marble runs [learningexplorationstrategiessolve].
F Embodied Interaction
World models provide the possibility for modeling virtual physical scenarios, yet a substantial gap persists between simulation environments and the real world. This gap highlights an essential limitation: modeling alone is insufficient for mastering the dynamics of real physical processes. Embodied intelligence addresses this limitation by requiring agents to perform multi-dimensional reasoning and interact with their physical surroundings. As a result, it moves beyond approaches that operate solely within controlled or virtual spaces and is increasingly viewed as a key paradigm for advancing future physics-based AI.
F.1 Robotics
While conventional world models excel at predicting outcomes in simulated environments, robotics applications demand the critical transition from passive observation to active manipulation. This evolution is exemplified by the emergence of Vision-Language-Action (VLA) models [VLA] that bridge the gap between building internal world representations and changing external physical reality. Pioneering systems like Gato [ReedEtAl2022_Gato] and RT-1/2 [BrohanEtAl2022_RT1, Brohan2023RT2VM] establishing the foundation for this paradigm shift.
Recent advances including $\pi_{0}$ [BlackEtAl2024_pi0], OpenVLA [KimEtAl2024_OpenVLA], and Gemini Robotics [DeepMind2024_GeminiRobotics] further demonstrate how physics-enhanced AI with embodiment extends beyond simulation-based world models by requiring real-time adaptation to physical constraints, sensor noise, and the irreversible consequences of actions in the real world. Figure 7 illustrates primary paradigms for current VLA systems. Both fusion-based (late/early) and dual-system architectures reflect the importance of integrating fast physical control with deliberative reasoning.
[…]
Figure 7: Architecture patterns for vision-language-action models. The shift from Late Fusion single-stream architectures to Dual-System designs reflects the integration of physical reasoning into VLAs.
The transition from simulated world models to reality systems also introduces new challenges. First, continuous action generation highlights the inadequacy of discretized representations when confronting the smooth, continuous nature of reality, because robotics demands continuous control that respects physical constraints [Safety-Critical]. Models like $\pi_{0}$ address this through flow matching [lipman2023_flow_matching], generating physically plausible trajectories that maintain stability during contact-rich manipulation. This direction is further explored through diffusion policies [ChiEtAl2023_DiffusionPolicy] . Furthermore, cross platform generalization tests whether learned physical principles can transcend specific hardware implementations. OpenVLA’s approach of learning normalized action deltas from diverse datasets [BrohanEtAl2022_OpenX] demonstrates how embodied model must abstract beyond simulation specific assumptions to achieve morphology independent reasoning that generalizes across varied physical platforms. Third, real-world perceptual grounding requires models to construct accurate world representations from noisy, partial observations. Some works addresses this through enhanced multi-view correspondence understanding [DeepMind2024_GeminiRobotics], object localization [LingEtAl2023_ActiveLearning, SeoEtAl2023_HybridControl] and physical properties recognition [WuEtAl2023_ObjectProperty], capabilities essential for reasoning about physical interactions under perceptual uncertainty.
F.2 Navigation
Navigation, a cornerstone of robotics, has evolved from a classical geometric problem to a core challenge in embodied intelligence, demanding interaction with and reasoning about the physical world. Current representative navigation tasks include object-goal navigation, vision-and-language navigation, dialog-based navigation, etc.
Object-Goal Navigation. Object-Goal Navigation (ObjectNav) tasks an agent to interact with its physical world by exploring an unknown space to find a target object, requiring it to reason about potential object locations based on visual cues. It’s best represented by the Habitat-Matterport3D (MP3D) [Matterport3D] benchmark, and benchmarks like AI2-THOR [AI2-THOR], Gibson [Gibson], RoboTHOR [RoboTHOR], HM3D-OVON [HM3D-OVON] belong to this task.
Vision-and-Language Navigation. Vision-and-Language Navigation (VLN) challenges an agent to reason about the connection between language and the physical world, interacting with its environment by translating natural language commands into a specific navigation path. For example, R2R [mattersim, Matterport3D], RxR [rxr], REVERIE [REVERIE] and TOUCHDOWN [TOUCHDOWN] are benchmarks for this task.
Dialog-based Navigation. Dialog-based Navigation allows an agent to interact with a user and its physical world simultaneously, reasoning about ambiguous instructions through conversation to successfully navigate its environment. Related to it are RobotSlang [RobotSlang], R2H [R2H], CVDN [CVDN], UNMuTe [UNMuTe] and so on.
Evolution of Navigation Methods. Navigation tasks have seen a significant methodological evolution, moving from early, specialized systems to models powered by large language models (LLMs). Early non-LLM-based methods were typically trained on domain-specific data to create a direct mapping from sensory inputs to navigation actions [zhang2024visionandlanguagenavigationtodaytomorrow]. This reliance on implicit representations often led to models that learned spurious correlations between inputs and outputs, limiting their generalization to unseen environments and making their decisions difficult to interpret [zhang2025researchnavigationmethodsbased].
The advent of LLMs and VLMs has introduced a new paradigm. Pre-trained on vast, diverse datasets, these models bring a wealth of real-world commonsense and physical knowledge that dramatically improves navigation precision. This shift has enabled explicit reasoning, such as CoT [CoT], which not only enhances decision accuracy but also provides greater interpretability [zhang2025researchnavigationmethodsbased]. This has led to a new wave of methods which are shown in Table IV and Figure 8.
Figure 8: Overview of NavCoT (Figure used courtesy of [NavCoT]).
TABLE IV: Representative methods for navigation tasks: Non-LLM vs. LLM-based
Task Method Non-LLM LLM-based
ObjectNav GOSE [chaplot2020objectgoalnavigationusing] RIM [RIM] LOAT [LOTA] LGR [LGR] CL-CoTNav [CL-CoTNav] ASCENT [ASCENT]
VLN BabyWalk [BabyWalk] ADAPT [ADAPT] HAMT [HAMT] EnvDrop [EnvDrop] DUET [DUET] NavGPT [NavGPT] NavCoT [NavCoT] VELMA [VELMA] NaVILA [NaVILA]
Dialog-based CMN [CMN] GVDN [GVDN] LLM- ContextBridge [ContextBridge] FLAME [FLAME]
However, these LLM-based approaches are not without their issues. They are susceptible to hallucinations, where incorrect physical reasoning can lead to failures, such as imagining a non-existent path through an obstacle [deng2025llmgoodpathplanner]. This fragility is further demonstrated by research showing that minor changes to a prompt can reduce a robot’s task success rate by nearly 20% [wu2025uncoveringfragilitytrustworthyllms]. A fundamental limitation is the representational bottleneck caused by converting rich physical world data into a token format, which can result in a significant loss of detail and the inability to reason about low-level preconditions for actions [RIM, chaplot2020objectgoalnavigationusing]. This is addressed by integrating world models, which can mentally simulate action outcomes, providing a crucial, grounded, and predictive layer that LLMs inherently lack [WebAgents]. This paradigm is also essential for bridging the "sim-to-real" gap, as a world model can learn a robust predictive representation of dynamics in simulation and transfer this knowledge to a real-world agent, as demonstrated by the TWIST [TWIST] framework’s teacher-student distillation approach.
F.3 Autonomous Driving
Autonomous driving refers to a system’s capability to perform part or all of the dynamic driving task (DDT) on a sustained basis without direct human intervention. The DDT includes all of the real-time operational and tactical functions required to operate a vehicle, such as steering, acceleration, and braking. To achieve this, an autonomous vehicle’s closed-loop software system integrates a series of core functions: perception, prediction, planning, and control [thorn2018framework]. Perception involves collecting and processing real-time data from various sensors like cameras and LiDAR to detect objects and road conditions [Grigorescu_2019]. Prediction forecasts the behavior of other road users [Rudenko_2020]. Planning generates a safe and efficient path for the vehicle to follow. Control translates the planned path into physical commands for the vehicle’s actuators [paden2016surveymotionplanningcontrol]. These tasks are implemented through several architectural paradigms [bojarski2016endendlearningselfdriving]. Current autonomous driving technology methods mainly include: Rule-based Approaches, Learning-based Approaches, World Models and Generative Approaches, Hybrid Approaches, etc.
Rule-based Approaches. Rule-based approaches rely on deterministic algorithms and explicit domain knowledge, encoding physics and traffic rules in interpretable models. Classical methods include the Intelligent Driver Model (IDM) [IDM] for car-following, RRT/RRT* [lavalle1998rapidly] for sampling-based trajectory generation, Model Predictive Control (MPC) [MPC] for optimization-based planning, and potential-field navigation [khatib1986real]. They are transparent and safe but struggle in dense and uncertain traffic. Benchmarks include CommonRoad [CommonRoad], nuPlan [nuplan], nuScenes [nuscenes], Waymo Open Dataset (WOD) [Waymo], CARLA [CARLA], and AirSim [AirSim].
Learning-based Approaches. Learning-based approaches replace rules with neural policies trained from data. Imitation Learning (IL) follows expert trajectories (e.g., ChauffeurNet [ChauffeurNet]), Inverse Reinforcement Learning (IRL) recovers reward functions from demonstrations [Maximum], and Reinforcement Learning (RL) optimizes policies in simulators [kiran2021deepreinforcementlearningautonomous]. AlphaDrive [AlphaDrive] and AutoDrive-R2 [AutoDrive] extend RL-based training with physics-informed rewards, reinforcing the role of simulation and safety-oriented benchmarks in this category. More recent methods use transformer-based forecasting (e.g., LaneGCN [LaneGCN], Wayformer [Wayformer]).
Generative Approaches. World models learn latent spatio-temporal representations of the driving environment, enabling rollouts of future states [World_models2018]. The Dreamer family (DriveDreamer [DriveDreamer] in Figure 9 and ReconDreamer [ReconDreamer]) capture compact latent dynamics for long-horizon simulation. Diffusion-based models include DrivingDiffusion [DrivingDiffusion], BEVControl [BEVControl], and Vista [Vista]. Transformer-based models such as GAIA-1/2 [GAIA-1] supports multimodal controllable scene generation. Occupancy-based models predict 4D voxelized space, such as OccWorld [OccWorld], RenderWorld [RenderWorld] and DOME [DOME] combine occupancy forecasting with planning via diffusion. Point-cloud generators like LiDARGen [LiDARGen] and the unified HERMES [HERMES] framework produce future LiDAR sweeps or merged 3D representations. Benchmarks in this domain include prediction and occupancy forecasting benchmarks (Occ3D-nuScenes/Waymo [Occ3D], OpenScene [OpenScene], Precog [PRECOG]), simulation benchmarks (CarlaSC [MotionSC], Bench2Drive [Bench2Drive]), and end-to-end planning benchmarks (nuPlan [nuplan], S2R-Bench [S2R-Bench]). These collectively test world models across scene generation, motion forecasting, and driving policy integration.
Figure 9: Overall framework of DriveDreamer (Figure used courtesy of [DriveDreamer]).
Hybrid Approaches. Hybrid approaches combine rule-based reliability with learning-based adaptability and cognitive reasoning. Neural waypoint predictors refined by MPC or rule-based controllers augmented with learned forecasting exemplify this paradigm [Gupta_2023]. Cognitive reasoning extensions include DriveCoT [DriveCoT], which integrates Chain-of-Thought supervision into driving tasks, and PRIMEDrive-CoT [PRIMEDrive-CoT], which introduces uncertainty-aware reasoning. PlanAgent separates high-level planning from execution, while LeapVAD [LeapVAD] adopts a dual-process architecture inspired by human cognition. DriveLMM-o1 [DriveLMM-o1] and Reason2Drive [Reason2Drive] further contribute cognition-augmented datasets for training and evaluating reasoning.
The current challenges in autonomous driving are fundamentally tied to the "long tail" problem of rare, high-impact edge cases that are difficult to encounter and address through traditional real-world data collection [wang2025terasimworldworldwidesafetycriticaldata, ren2025cosmosdrivedreamsscalablesyntheticdriving]. World models solve this by creating a learned, internal simulation of the environment, which functions as a "computational snow globe" where the AI can "mentally rehearse" actions [Think2Drive, hafner2024masteringdiversedomainsworld]. By generating and training in an unbounded number of synthetic, safety-critical scenarios, these models allow the autonomous system to anticipate and mitigate hazardous situations before they unfold in the physical world, ultimately shifting the paradigm from reactive to predictive autonomy and improving real-world safety [ding2023surveysafetycriticaldrivingscenario].
G Discussion
The Isolation of Perception and Reasoning. This survey reveals a fundamental paradox: AI systems achieve superhuman performance on isolated physical tasks, e.g. olympiad-level problem solving, photorealistic video generation, yet lack the flexible, principle-based understanding that allows a child to predict whether stacked blocks will topple. This gap exposes not a quantitative shortfall in data or compute, but a qualitative misalignment between how current systems learn physics and how physical understanding actually works. The four capabilities examined (physical perception, physics reasoning, world modeling and embodied interaction) are not separate research directions but facets of a unified cognitive architecture that current approaches fail to integrate. Physical perception without reasoning remains trapped in correlation; physics reasoning without perceptual grounding produces detached symbolic manipulation; world models without embodied feedback generate plausible but physically inconsistent predictions. Therefore, we advocate that the research community fundamentally rethink how models can mutually inform and constrain these four capabilities through bidirectional coupling, rather than pursuing isolated task improvements.
Sim-to-real Gap. The transition from simulation to embodied interaction exposes the field’s most critical vulnerability. World models trained on internet-scale video generate visually compelling predictions while systematically violating physical principles, e.g., objects floating without support, collisions without momentum transfer. This fragility stems from optimizing for perceptual plausibility rather than physical consistency. Without embodied consequences to enforce correct physical behavior, systems learn only superficial correlations. We should place greater emphasis on current world models’ simulation-to-reality transfer capabilities, for example, by actively refining physical models based on prediction errors encountered during interactions, treating real-world feedback as an essential component of learning rather than merely an end-to-end deployment task.
Internalizing Natural Laws. Current approaches reveal a deeper architectural limitation: most systems, including frontier multimodal language models, rely on pattern matching over vast datasets rather than internalizing the compositional, causal structure of the law of physics. This explains why models achieve 50-80% accuracy on in-distribution benchmarks yet fail catastrophically on counterfactual scenarios or novel configurations. Physics is not a collection of statistical regularities but principles like conservation laws, symmetries, causal mechanisms that compose systematically. The path forward demands architectures that encode physical law to overcome inductive bias: differentiable physics engines maintaining hard constraints, neuro symbolic systems integrating learned perception with symbolic reasoning, or embodied learning paradigms acquiring intuition through active intervention rather than passive observation. These architectural innovations provide the possibility for internalizing natural laws. We have reason to believe that, beyond the scaling laws of data, parameter and inference time, a promising path forward may lie in scaling laws grounded in the rules of physical world.
H Conclusion
This survey presents a task-oriented taxonomy tracing physical AI development through four progressive capabilities: physical perception, physics reasoning, world modeling, and embodied interaction. Each stage builds upon the previous, advancing from passive observation to active physical comprehension. Despite progress in individual domains, most systems still rely on statistical pattern recognition rather than genuine physical understanding. Moving forward, the most promising path lies in hybrid approaches that integrate physics-grounded architectures, physics-informed training, and symbolic reasoning into unified frameworks. Such frameworks that traverse the entire progression from perception to interaction will enable AI systems to genuinely perceive, reason about, model, and interact with physical reality for safe and reliable deployment.